A farmer has a rectangular garden plot surrounded by 200 ft of fence. Find the length and width of the garden if its area is 2400 square feet.
Answer by nerdybill(7003)

A farmer has a rectangular garden plot surrounded by 200 ft of fence. Find the length and width of the garden if its area is 2400 square feet.
.
Let x = width
and y = length
.
The because he has 200 ft of fencing, this must lest than or equal to the perimeter:
2(x+y) = 200
x+y = 100
Solving for x:
x+y = 100
x = 100-y
.
Area:
xy = 2400
.
substitute x=100-y:
xy = 2400
(100-y)y = 2400
100y-y^2 = 2400
0 = y^2-100+2400
factoring:
0 = (y-40)(y-60)
y={60,40}
.
length 60 feet
width 40 feet